OpenStudy (anonymous):

How to find the inverse Laplace transform on Ti-89?

OpenStudy (anonymous):

e TI-89 has a function Laplace (f(t),t) which will give the Laplace transform of a function. The inverse Laplace transform can be found with Laplace(f(t),t) You can find these functions in the calculator's catalog.

OpenStudy (anonymous):

But when I click catalog, I couldn't find it.

OpenStudy (anonymous):

That's not what I'm saying. I want to know how to do it both in calculator and by doing the work.

OpenStudy (loser66):

why? Laplace transform and inverse Laplace transform table are available and we are allowed to use it. what 's your problem, let see what can I do to help

OpenStudy (anonymous):

Even on the test, we're allowed to look at the table? Is there a way that we can solve the problems without looking at the table?

OpenStudy (loser66):

sure, You should ask your prof about that. the table is the key to solve problem. we use it back and forth. cannot remember all of them.

OpenStudy (loser66):

however, even though you have table, it's just a small step. Have to solve it by hand after that small step.

OpenStudy (anonymous):

Okay, I was working on this problem. Find the inverse Laplace transform of F(s)=(2s-3)/(s^2-4).

OpenStudy (loser66):

Let me send you my table.

OpenStudy (loser66):

I will scan and post the answer, you just follow. I know this field is not easy to understand.

OpenStudy (anonymous):

Now I get part of the answer with cosh 2t and sinh 2t. But how can I find the 2 roots?

OpenStudy (loser66):

that's it.

OpenStudy (anonymous):

Yes, I got it. Thanks a lot. The table really helped. Are you a college junior or senior?

OpenStudy (anonymous):

I'm a college junior.

OpenStudy (anonymous):

I don't have a major.

OpenStudy (anonymous):

I'm about to drop out of college. The professors are so f*cking weird. I hate them.

OpenStudy (loser66):

my prof is from Russia. I cannot understand what he talks. just self-study

OpenStudy (anonymous):

It must be tough then.

OpenStudy (loser66):

why do you have to take this course?

OpenStudy (anonymous):

I don't have to. I wanted to study higher math.

OpenStudy (loser66):

ok, good boy.

OpenStudy (loser66):

ah, you need roots. sorry, I didn't understand what you meant. Now I got it. you have \[2 cosh( 2t) -\frac{3}{2} sinh (2t) =2\frac{e^{2t}+e^{-2t}}{2}-\frac{3}{2}\frac{e^{2t}-e^{-2t}}{2}\] \[=\frac{1}{4}\left(e^{2t}+7e^{-2t}\right)\] ahahaha.. my answer is exactly what Wolframalpha calculate. take a look https://www.wolframalpha.com/input/?i=inverse+laplace+transfom+%282s-3%29%2F%28s^2-4%29 the last box. lalala... life is beautiful